OpenStudy (anonymous):

help me with this problem please .. 15n^2+19n-10=0

OpenStudy (lalaly):

(3n+5)(5n-2)

OpenStudy (lalaly):

=0 n=-5/3 , n=2/5

OpenStudy (anonymous):

thank you so much .. last 1 problem haha

OpenStudy (lalaly):

lolz no problm:)